OpenStudy (anonymous):

Which of the following equations does not represent the line in the graph shown above? A. y = −(1/2)x + 2 B. y = (1/2)x + 2 C. y − 2 = −(1/2)x D. (1/2)x = 2 − y

OpenStudy (mathstudent55):

Let's rewrite C. and D. so they are in the y = mx + b form. C. y - 2 = -(1/2)x y = -(1/2)x + 2 D. (1/2)x = 2 - y y + (1/2)x = 2 y = -(1/2)x + 2 Now let's look at all the choices again in y = mx + b form: A. y = −(1/2)x + 2 B. y = (1/2)x + 2 C. y = -(1/2)x + 2 D. y = -(1/2)x + 2 Notice that three of the choices are the same equation. Only one choice is a different equation. Also, notice that the three equations that are the same have a negative slope. That means this is a line that tilts down to the right. The equation that is different has a positive slope which means a line that tilts up to the right.

OpenStudy (anonymous):

So it would be B